Another pre-test question. I can't seem to figure out how to start on these things.
Solve.
x^2=-6x
Do I subtract the six like so -
x^2+6=0

Answer by Nate(3500)

x^2 = -6x
x^2 + 6x = 0
x(x + 6) = 0
x = 0 and x + 6 = 0
x = 0 and x = -6